OpenStudy (scarlettfarra2000):

Which method is most efficient method to use to solve 2x^2+5x-11=0

OpenStudy (solomonzelman):

The discriminant is: \(\color{#000000}{ \displaystyle \Delta = b^2-4ac }\) \(\color{#000000}{ \displaystyle a=2 }\) \(\color{#000000}{ \displaystyle b=5 }\) \(\color{#000000}{ \displaystyle c=-11 }\) In this case, \(\color{#000000}{ \displaystyle \Delta = 5^2-4(2)(-11)=25+88=113 }\)

OpenStudy (solomonzelman):

Since the discriminant is not a perfect square, therefore you can not factor this equation.

OpenStudy (solomonzelman):

So, I would use the quadratic formula.

OpenStudy (scarlettfarra2000):

Ok, thanks just making sure

OpenStudy (solomonzelman):

No Problem.